# DERD-Net: Learning Depth from Event-based Ray Densities

### v1

@@ 4.3 Experiments on MVSEC Dataset
Algorithm Modality Mean Err Median Err bad-pix #Points [cm] [cm] [%] [million] SOTA EMVS [31] monocular + EMVS [31] monocular + ESVO [46] stereo ESVO indep. 1s stereo SGM indep. 1s stereo GTS indep. 1s stereo MC-EMVS [13] stereo + MC-EMVS [13] stereo + MC-EMVS [13] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +
| Scene | Pixel selection map | MC-EMVS [13] | MC-EMVS + Fdenser | DERD-Net (Ours) | Ground truth (GT) |
Algorithm Modality Mean Err Median Err bad-pix #Points [m] [m] [%] [million] SOTA EMVS [31] monocular + EMVS [31] monocular + ESVO [46] stereo MC-EMVS [13] stereo + MC-EMVS [13] stereo + MC-EMVS [13] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +
In this section, we describe the experiments conducted on the indoor_flying sequences 1,2,3 of the MVSEC dataset [48] to evaluate the performance of the proposed depth estimation method. We employed three-fold cross-validation by utilizing two sequences for supervised training and reserving the remaining sequence for testing, repeating this process for all three possible combinations of sequences to ensure robustness in our evaluation.
@@
Results. The averaged results over all three sequences are displayed in Tab. 3 for all discussed modalities, while the individual results (including performance without EL) are detailed in the Supplementary Material, in Tabs. 9, 10 and 11. Notably, our network’s performance converges after only 3 epochs of training. This rapid convergence is particularly advantageous for future applications where the network might be trained on more heterogeneous datasets or retrained for specific scenarios.
@@
Stereo setup. Applying our single-pixel network to stereo DSIs filtered by allows us to predict depth at significantly more pixels than any other method, except for the SGM method [17], while consistently surpassing all benchmarks across all metrics. The number of pixels increases by 242%, while the MAE and MedAE reduce by 24% and 30%, compared to MC-EMVS [13] with . The only exception is the bad-pix measure, where MC-EMVS performs slightly better. When both methods are compared on the same set of pixels, our approach yields a reduction in both MAE and MedAE of 42% for and 46% for , respectively. Performance remains consistent when using the multi-pixel network, which increases the amount of depth-estimated pixels by a factor of 5.47 for and 4.09 for compared to its single-pixel version, indicating it to be a superior approach to the morphological filter of MC-EMVS, which only rises the number of points by a factor of 3.70 for . As a consequence, the multi-pixel version of our framework estimates depth for almost as many pixels as the SGM method while delivering new SOTA results. Additionally, while no other learning-based methods report good generalization on indoor_flying_2 (see Tab. 10 in the Supplementary Material) due to differences in dynamic characteristics between training and testing events in that split [2, 39], our network excels, underscoring its robust generalization abilities.
Qualitative comparison. To further illustrate the effectiveness of our method, Fig. 3 compares depth maps generated by our single-pixel network against those produced by SOTA method MC-EMVS. Our network not only provides a denser depth estimation, which improves the recognition of contours, but also effectively eliminates the visible outliers produced by MC-EMVS. This improvement is evident when comparing our method to MC-EMVS applied both to the expanded and the original set of pixels, highlighting the robustness and superiority of our approach.
@@ 4.4 Experiments on DSEC Dataset
To assess the applicability of our network architecture to different data, we retrained and tested it on DSIs obtained from a stereo setting in the DSEC dataset [10]. This dataset presents unique challenges due to its outdoor driving scenarios, which differ significantly from the indoor environments of the MVSEC dataset, its higher spatial resolution (640 480 px) and different noise characteristics (Prophesee camera vs. DAVIS346 camera). Moreover, straight driving sequences are especially challenging for event-based multi-view stereo due to little motion parallax.
@@
We selected the Zurich_City_04_a sequence to provide a focused in-depth evaluation. We split the sequence into two halves for training and testing. The DSIs were created by fusing the left and right DSIs via the harmonic mean. Analogous to Sec. 4.3, we use the original filter from MC-EMVS [13] with a window size of px and , and a denser filter with a window size of px and . First, the network was trained for 3 epochs on the DSIs of the first half of the sequence and tested on those of the second. The process was then reversed and each network was used to predict in its testing half of the data sequence.
@@
The results of these experiments are displayed in Tab. 4 and illustrated in Fig. 3. Our approach drastically outperforms every other method across all metrics, with our multi-pixel network achieving even slightly better performance than the single-pixel network. For , it reduces the MAE by 55% on a 1.72x higher number of pixels compared to MC-EMVS with a morphological filter. For , depth estimation density is increased by an additional factor of 2.24 while performance remains mostly stable, yielding a reduction in MAE of 62% compared to the argmax operation from MC-EMVS. Remarkably, even on purely monocular DSIs filtered by , our framework achieved superior performance to all benchmarked methods for every metric except MedAE. These results underscore the robustness and versatility of our approach, even in complex real-world outdoor scenes.
Reflection Modality Mean Err Median Err bad-pix #Points [cm] [cm] [%] [million] none stereo + vertical stereo + horizontal stereo + horizontal + vertical stereo + none stereo + vertical stereo + horizontal stereo + horizontal + vertical stereo +
Sub-DSI Modality Mean Err Median Err bad-pix #Points frame size [cm] [cm] [%] [million] stereo + stereo + stereo + stereo + Fdenser stereo + Fdenser stereo + Fdenser
4.5 Data Augmentation Analysis
To further analyze the generalization capabilities of our network, we applied axis-aligned reflections to the DSIs generated from the indoor_flying_1 sequence of the MVSEC dataset. Specifically, we flipped the DSIs horizontally, vertically, and both horizontally and vertically. These transformations allow us to evaluate the network’s performance on scenes that share the same configuration (e.g., distance ranges) but differ due to changes in viewpoint. We retested the single-pixel network that was trained solely on the original indoor_flying 2 and 3 sequences, without retraining it or applying any form of data augmentation during the original training process. This ensures an unbiased evaluation of the network’s inherent ability to generalize to new scenes. Importantly, data augmentation could be used during training to further enhance performance, but for this study, we purposefully avoided it to provide a clear assessment of the network’s natural generalization capabilities.
The results of these experiments are displayed in Tab. 5. Performance only worsened slightly, with results that still significantly outperform all SOTA methods for all tested configurations, indicating that our network might effectively generalize to scenes that share similar depth ranges and texture to those on which it was originally trained.
4.6 Sensitivity Analysis
The hyperparameters used for our experiments are detailed in Tab. 2. The LiDAR interval, the estimated minimum and maximum distances, and the number of depth layers have all been defined to match the protocol in [13]. In Sec. 4.3, we already provided a comparison between the two AGT filters and . In this subsection, we therefore analyze the impact of the size of the Sub-DSI.
We retrained the network for one epoch on the indoor_flying 2 and 3 sequences and compared its test performance on sequence 1 using a radius of 2, 3, and 4 px, effectively creating Sub-DSI frames of size , , and px. From the results reported in Tab. 6 it can be inferred that performance appears to improve as the radius increases. While data size increases proportionally as well, the network was able to achieve better results after a single epoch using a frame size of than when fully trained on frames (see Tab. 9 in the Supplementary Material), highlighting its potential for further performance improvements.
@@ 4.7 Runtime
To assess the efficiency of our network, we measured its inference time. It achieved an average of only 0.37 ms per Sub-DSI on an NVIDIA RTX A6000. Since predictions are made independently per pixel, inference for each Sub-DSI can be parallelized on the GPU. The total inference time to estimate a depth map of average density (500 pixels) from MVSEC with is 1.12 ms. This ultra-fast performance, combined with our network’s lightweight architecture, enables efficient execution, making it ideal for real-world applications requiring low-latency depth estimation.
4.8 On robustness compared to other deep-learning–based depth estimation methods
Previous sections have compared our approach to depth estimation methods that have the same input (events and camera poses) and output (semi-dense depth maps), i.e., that tackle the same task (depth prediction for SLAM). This section also provides Tab. 7 to compare to end-to-end learning-based stereo methods, which are “instantaneous” (do not take into account camera poses) and output dense depth. Applying these methods for tracking would therefore require an extra step of depth-pixel selection to avoid becoming computationally expensive.
Since errors come from different observation windows and ground truth points, absolute accuracy is not directly comparable [14]. Nevertheless, comparing the errors in the different splits relative to each other is informative about robustness: we observe that our method is the first one to generalize robustly across all three sequences. No other method reports good generalization on “split 2” of MVSEC because of the difference in dynamic characteristics of events in training and testing on that split (as mentioned in [38, 1]). This is true even when compared to hybrid approaches, despite them also using stereo intensity frames (“2E+2F” input data modality).
5 Discussion and Conclusion
@@
We have developed the first learning-based multi-view stereo method for event-based depth estimation. Our approach combines input camera poses with events to produce intermediate geometric representations (DSIs) from which depth is estimated using deep learning. It is directly applicable to both monocular and stereo camera setups. By processing small independent subregions of DSIs in parallel, the framework operates independently of camera resolution and facilitates an efficient network under 1 MB in size with an inference time of only 0.37 ms. Its constant low model complexity and memory cost, regardless of the event rate, provide full scalability and enables efficient execution.
@@
Our framework consistently demonstrated superior performance across several metrics compared to other stereo methods and achieved comparable performance when using purely monocular data. It is the first learning-based depth estimation approach that reports robust generalization on all three indoor flying sequences of the MVSEC dataset. Adaptability to different scenes was confirmed on the outdoor driving DSEC dataset, for which it drastically outperformed benchmark approaches across all metrics. Moreover, our framework significantly increased the number of points for which depth can be robustly estimated from DSIs.
@@
Given its exceptional performance, ultra-lightweight architecture, scalability and flexibility across different configurations, our method holds strong potential to become a standard approach for learning depth from events and is highly suitable for real-world applications such as SLAM.
@@ Acknowledgments
This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135.
Appendix A Supplementary Material

### v2

@@ 4.3 Experiments on MVSEC Dataset
| Scene | Pixel selection map | MC-EMVS [12] | MC-EMVS [12] with Fdenser | DERD-Net (Ours) | Ground truth (GT) |
This section describes the experiments conducted on the indoor_flying sequences 1,2,3 of the MVSEC dataset [4] to evaluate the performance of the proposed depth estimation method. Most stereo methods do not evaluate on indoor_flying_4 (because of noisy events from the low-texture floor as the drone flies very low) and the driving sequences (because the stereo baseline is too small for the given depth range and low camera resolution). We employed three-fold cross-validation by utilizing two sequences for supervised training and reserving the remaining sequence for testing, repeating this process for all three possible combinations of sequences to ensure robustness in our evaluation.
@@
Results. The averaged results over all three sequences are displayed in the left half of Tab.˜3 for all discussed modalities, while the individual results (including performance without EL) are detailed in the Appendix, in Tabs.˜12, 13 and 14. Notably, our network’s performance converges after only 3 epochs of training. This rapid convergence is particularly advantageous for future applications where the network might be trained on more heterogeneous datasets or retrained for specific scenarios.
@@
Stereo setup. Applying our single-pixel network to stereo DSIs filtered by allows us to predict depth at significantly more pixels than any other method, except for the SGM method [38], while consistently surpassing all benchmarks across all metrics. The number of pixels increases by 242%, while the MAE and MedAE reduce by 24% and 30%, compared to MC-EMVS [12] with . The only exception is the bad-pix measure, where MC-EMVS performs slightly better. When both methods are compared on the same set of pixels, our approach yields a reduction in both MAE and MedAE of 42% for and 46% for , respectively. Performance remains consistent when using the multi-pixel network, which increases the amount of depth-estimated pixels by a factor of 5.47 for and 4.09 for compared to its single-pixel version, indicating it to be a superior approach to the morphological filter of MC-EMVS, which only rises the number of points by a factor of 3.70 for . As a consequence, the multi-pixel version of our framework estimates depth for almost as many pixels as the SGM method while delivering new SOTA results.
Qualitative comparison. To further illustrate the effectiveness of our method, Fig.˜3 compares depth maps generated by our single-pixel network against those produced by SOTA method MC-EMVS. Our network not only provides a denser depth estimation, which improves the recognition of contours, but also effectively eliminates the visible outliers produced by MC-EMVS. This improvement is evident when comparing our method to MC-EMVS applied both to the expanded and the original set of pixels, highlighting the robustness and superiority of our approach.
@@ 4.4 Experiments on DSEC Dataset
To assess the applicability of our network architecture to different data, we retrained and tested it on DSIs obtained from a stereo setting in the DSEC dataset [13]. This dataset presents unique challenges due to its outdoor driving scenarios, which differ significantly from the indoor environments of the MVSEC dataset, its higher spatial resolution (640 480 px) and different noise characteristics (Prophesee camera vs. DAVIS346 camera). Moreover, straight driving sequences are especially challenging for event-based multi-view stereo due to the little motion parallax present in them.
@@
Setup. We select the commonly used Zurich_City_04_a sequence to provide a focused in-depth evaluation. We split the sequence into two halves for training and testing. The DSIs were created by fusing the left and right DSIs via the harmonic mean. Analogous to Sec.˜4.3, we use the original filter from MC-EMVS [12] with a window size of px and , and a denser filter with a window size of px and . First, the network was trained for 3 epochs on the DSIs of the first half of the sequence and tested on those of the second. The process was then reversed and each network was used to predict in its testing half of the data sequence.
@@
Results. The results of these experiments are displayed on the right half of Tab.˜3 and illustrated in Fig.˜3. Our approach drastically outperforms every other method across all metrics, with our multi-pixel network achieving even slightly better performance than the single-pixel network. For , it reduces the MAE by 55% on a 1.72x higher number of pixels compared to MC-EMVS with a morphological filter. For , depth estimation density is increased by an additional factor of 2.24 while performance remains mostly stable, yielding a reduction in MAE of 62% compared to the argmax operation from MC-EMVS. Remarkably, even on purely monocular DSIs filtered by , our framework achieved superior performance to all benchmarked methods for every metric except MedAE. These results underscore the robustness and versatility of our approach, even in complex real-world outdoor scenes.
4.5 Robustness of DERD-Net compared to other deep-learning stereo methods
Since there are no comparable learning-based methods that use prior camera poses, Tab.˜4 compares end-to-end learning-based stereo methods, which are “instantaneous” (do not take into account camera poses) and output dense depth. In order to use their output for efficient VO/SLAM, we would need an extra step of extracting features (keypoints). Instead, DERD-Net’s semi-dense depth maps help avoid unnecessary computation by outputting 3D edges for direct visual odometry, as in [14].
We use the same train-test splits established as the other learning-based methods [5]. While absolute accuracy is not directly comparable, evaluating the errors in the different splits relative to each other is informative about robustness: we observe that our method is the first one to generalize robustly across all three sequences (Tab.˜4). No other method reports good generalization on “split 2” of MVSEC because of the difference in dynamic characteristics of events in training and testing on that split [6, 43]. This is true even when compared to hybrid approaches, despite them also using stereo intensity frames (“2E+2F” input data modality). The observed robustness of our method to such shifts may be supported by the architectural choice of processing only small subregions as input (see Tab.˜2 for Sub-DSI frame size), which encourages the model to learn generalizable patterns within the Sub-DSIs rather than memorizing global scene layout or dataset-specific context.
4.6 Sensitivity Analyses
In this section we carry out experiments varying the settings in Tab.˜2. Furthermore, we analyze the robustness of our method to noisy camera poses obtained from an event-based SLAM system.
Sensitivity with respect to sub-DSI size. Varying the horizontal and vertical extent of the Sub-DSIs has an impact on our method’s performance. Our experiments show that the performance of DERD-Net can be improved by increasing the frame size of the Sub-DSIs, at the expense of increasing the network complexity (e.g., parameter count and computational cost). See Appendix Sec.˜A.1.
Sensitivity with respect to DSI transformations. We analyzed how DERD-Net behaves in the case of previously unseen but structurally similar environments, obtained by means of horizontal and vertical flips of the DSIs. Although its performance worsened slightly, it still outperformed all baseline methods. This demonstrates robustness to the aforementioned transformations. See Appendix Sec.˜A.2.
Sensitivity with respect to noisy camera poses. To assess the importance of having accurate camera poses during DSI construction, we test our framework using noisy poses with drift, mimicking real-world SLAM conditions. Instead of ground-truth (GT) poses from LiDAR-IMU odometry, we use poses estimated by the stereo event-based VO system ES-PTAM [14], which reports an Absolute Trajectory Error (ATE) of 131.62 cm over a 50 m-deep scene in the DSEC Zurich_City_04_a sequence. Running DERD-Net with these imperfect poses yields the results shown in the top rows of Tab.˜5. The percentage values in parentheses denote the relative differences with respect to the performance obtained using ideal (GT) poses (Tab.˜3). Remarkably, performance improved across all metrics (likely due to the slight reduction in the number of evaluated points of comparable magnitude), demonstrating strong robustness of DERD-Net to noisy poses obtained from an event-based SLAM system.
We conduct an additional experiment where the original DERD-Net depth predictions were used to re-estimate the camera poses (in an offline manner, using the camera tracking module in [14]). The resulting poses were then used to build DSIs on which DERD-Net was evaluated. This “reprojection” loop allows us to assess, using standard depth-based metrics, the robustness of our method to noise in camera poses introduced by DERD-Net’s own depth inaccuracies. The results are reported in the bottom rows of Tab.˜5. The performance shows only minor degradation, particularly for , with no metric worsening by more than 13%. Remarkably, even under such self-induced pose noise, DERD-Net’s depth estimation errors remain roughly 50% lower than those of prior SOTA methods using ideal poses. These results demonstrate the practical viability of deploying DERD-Net as a depth-estimation module within a self-sustaining SLAM system. Overall, our experiments confirm that DERD-Net remains remarkably robust even when the input poses are significantly degraded, as would be expected in real-world scenarios. See also Appendix Secs.˜A.3 and A.4
Algorithm Poses Filter Mean Err Median Err bad-pix #Points [m] [m] [%] [million] DERD-Net ES-PTAM 1.56 0.45 3.84 1.61 (-3.11%) (-2.17%) (-6.8%) (-3.59%) DERD-Net ES-PTAM 1.74 0.52 4.84 4.49 (-3.33%) (-3.7%) (-3.97%) (-3.23%) DERD-Net Reprojection 1.66 0.49 4.19 1.46 (+3.11%) (+6.52%) (+1.7%) (-12.57%) DERD-Net Reprojection 1.95 0.60 5.65 4.09 (+8.33%) (+11.11%) (+12.1%) (-11.85%)
@@ 4.7 Runtime
Our network achieved an average inference time of only 0.37 ms per Sub-DSI on an NVIDIA RTX A6000. Since predictions are made independently per pixel, inference for each Sub-DSI can be parallelized on the GPU. The total inference time to estimate a depth map of average density (500 pixels) from MVSEC with is 1.12 ms.
Taking MVSEC as an example (DAVIS cameras of 346260 pixels) and DSIs back-projecting 2 million events onto depth planes, then each DSI creation takes 45 ms, DSI fusion takes 26 ms, and pixel selection takes 0.2 ms on an 8-core computer with Intel Xeon(R) W-2225 CPU operating at 4.10 GHz. These values are common for both the state-of-the-art method MC-EMVS and DERD-Net. It has been shown that DSI creation does not hamper real-time performance [53] because the 3D map can be updated infrequently and on-demand.
Our network adds only a very small runtime compared to the DSI creation time. This ultra-fast performance, combined with its lightweight architecture, enables efficient execution, making DERD-Net ideal for real-world applications requiring low-latency depth estimation.
5 Conclusion
@@
We have developed the first learning-based multi-view stereo method for event-based depth estimation. Our approach combines input camera poses with events to produce intermediate geometric representations (DSIs) from which depth is estimated using deep learning. It is directly applicable to both monocular and stereo camera setups. By processing small independent subregions of DSIs in parallel, the framework operates independently of camera resolution and facilitates an efficient network under 1 MB in size with an inference time of only 0.37 ms.
@@
Our framework consistently demonstrated superior performance across several metrics compared to other stereo methods and achieved comparable performance when using purely monocular data. It is the first learning-based depth estimation approach that reports robust generalization on all three indoor flying sequences of the MVSEC dataset. Adaptability to different scenes was confirmed on the outdoor driving DSEC dataset, for which it drastically outperformed benchmark approaches across all metrics. Moreover, our framework significantly increased the number of points for which depth can be robustly estimated from DSIs. It also showed strong robustness to noise in camera poses.
@@
Given its exceptional performance, ultra-lightweight architecture, scalability and flexibility across different configurations, our method holds strong potential to become a standard approach for learning depth from events and is highly suitable for real-world robotic applications requiring low latency and low memory such as SLAM [15].
@@ Acknowledgments
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135.
Appendix A Appendix: Sensitivity Analyses
In this section we report additional experiments to assess our method’s robustness to changes in hyperparameter (Tab.˜6), changes in the scene (Tab.˜7) and noise in camera poses (Tabs.˜8 and 9). We also complement the evaluation on the downstream task of camera tracking (Tab.˜10).
A.1 Sensitivity with respect to Sub-DSI Size
The hyperparameters used for our experiments are detailed in Tab.˜2. The LiDAR’s sampling interval, the estimated minimum and maximum depths, and the number of depth layers have all been defined to match the protocol in [12]. In Sec.˜4.3, we already provided a comparison between the two AGT filters and . In this section, we therefore analyze the impact of the size of the Sub-DSI.
We retrained the network for one epoch on the indoor_flying 2 and 3 sequences and compared its test performance on sequence 1 using radii of 2, 3, and 4 px, effectively creating Sub-DSI frames of size , , and px, respectively. Layer dimensions were adapted, while the overall network architecture remained fixed.
From the results reported in Tab.˜6 it can be inferred that performance appears to improve as the radii increase. The network was able to achieve better results after a single epoch using a frame size of than when fully trained on frames (see Tab.˜12 in this Appendix), highlighting its potential for further performance improvements.
Nevertheless, increasing the frame size to yielded a reduction of 5% in MAE for both filters after a single training epoch. In contrast to that, the network had to apply 65% more 3D-convolutional operations and its total amount of parameters raised from 70k to 270k. We therefore decided for a frame size for this study. Future research could explore the evident potential to further boost performance by optimizing the sub-DSI size, considering the trade-off between accuracy, parameter count and computational costs.
A.2 Sensitivity with respect to DSI Transformations
Next, we analyze the robustness of DERD-Net with respect to transformations of the DSI, in particular axis-aligned reflections of the DSIs generated from the indoor_flying_1 sequence of the MVSEC dataset. Specifically, we flipped the DSIs horizontally, vertically, and both horizontally and vertically. These transformations effectively generate scenes with similar geometric properties (e.g., distance ranges) but novel spatial configurations. This allows us to evaluate how well the network generalizes to previously unseen, yet structurally similar environments. We therefore purposely used no data augmentation during training to ensure a representative assessment of the network’s inherent robustness. Analogous to previous experiments, we used the single-pixel network that was trained solely on the original indoor_flying 2 and 3 for evaluation. No retraining was performed.
The results of these experiments are displayed in Tab.˜7. Performance worsened only slightly, with results that still significantly outperform all SOTA methods for all tested configurations, indicating that our network might effectively generalize to scenes that share similar depth ranges and texture with those on which it was originally trained.
A.3 Sensitivity with respect to Noise in Camera Poses
In Section Sec.˜4.6 we summarized the sensitivity of DERD-Net with respect to noise in the camera poses used to build DSIs, on DSEC data. For completeness, we now show results on MVSEC data.
We repeat the same experiment as that in the top rows of Tab.˜5 on the MVSEC indoor_flying_1 sequence, for which ES-PTAM reported an ATE of 14.93 cm over a 6 m depth range. The results shown in Tab.˜8, and compared to those obtained with GT poses in Tab.˜12, again highlight DERD-Net’s strong robustness to noisy poses estimated by an event-based SLAM system: the MAE and MedAE increased only slightly, while the bad-pix metric even improved. The most pronounced decline was in the number of evaluated points. Nevertheless, DERD-Net with still predicts depth for 69% more pixels than MC-EMVS with under ideal poses, while achieving a 30% lower MAE. For its multi-pixel variant, DERD-Net evaluated with noisy poses maintains superior performance over all state-of-the-art methods using GT poses, while still predicting the highest number of points.
Algorithm Poses Filter Mean Err Median Err bad-pix #Points [m] [m] [%] [million] DERD-Net ES-PTAM 12.72 6.33 0.65 0.68 (+8.81%) (+16.79%) (-24.42%) (-30.61%) DERD-Net ES-PTAM 15.76 7.36 1.17 1.62 (+6.06%) (+13.76%) (-10.69%) (-46.18%) DERD-Net (multi-pixel) ES-PTAM 13.53 6.76 0.65 3.41 (+10.27%) (+19.01%) (-24.42%) (-33.91%) DERD-Net (multi-pixel) ES-PTAM 16.60 7.65 1.21 6.27 (+6.62%) (+16.08%) (-11.68%) (-46.46%)
In the interest of thoroughness, we also used poses from the state-of-the-art event-based stereo visual–inertial odometry system ESVO2 [42], which is notably more accurate than ES-PTAM, to run DERD-Net on MVSEC indoor_flying_1, indoor_flying_2, and indoor_flying_3. The mean results are reported in Tab.˜9. Compared to Tab.˜3, DERD-Net shows only a slight decrease in depth estimation performance on , while it even improves on , confirming its robustness to pose noise from an event-based SLAM system integrating events and inertial data.
Algorithm Poses Filter Mean Err Median Err bad-pix #Points [m] [m] [%] [million] DERD-Net ESVO2 11.53 5.73 1.09 0.61 (-1.37%) (+4.18%) (+22.47%) (-22.78%) DERD-Net ESVO2 13.69 6.32 1.46 1.45 (-10.17%) (-5.39%) (-14.12%) (-47.65%)
A.4 Downstream Camera Tracking Performance Analysis
As intermediate results to those in the bottom part of Tab.˜5, we report offline camera tracking performance using the edge-alignment camera tracking module in [14] acting on input events and the local maps built using DERD-Net’s depth predictions (from GT poses). Camera tracking performance is given in terms of ATE and Absolute Rotation Error (ARE) on the DSEC [13] driving dataset in Tab.˜10. Although our model was trained only on the Zurich_City_04_a sequence, we evaluate it on all Zurich_City_04 sequences to highlight its generalization capabilities.
Sequence zc04a zc04b zc04c zc04d zc04e zc04f Duration [s] 35 13.4 53 47.8 13.6 43.1 ATE RMSE [cm] 17.07 7.85 14.00 55.64 5.71 36.11 ARE RMSE [deg] 0.31 0.08 0.45 0.67 0.11 0.72
The obtained pose errors in the 50 m depth range scenes across all sequences show strong performance of DERD-Net for downstream tasks such as pose estimation via simple photometric edge alignment on event images, as well as its robust generalization even when trained on a single sequence. Training on a more diverse set of DSEC sequences would be expected to further enhance these results. These values are not comparable to online SLAM tracking results because they assume that the 3D map was pre-built offline using DERD-Net with GT poses. Therefore, the estimated camera poses reported here do not accumulate drift.
Appendix B Appendix - Detailed per-Sequence Results